OpenStudy (anonymous):

Simplify the rational expression. State any excluded values. x-2/x^2+3x-10 A. x^2 B. x+5, where x is not equal to -5 C. 1/x2 where x is not equal to -2 D. 1/x+5 where x is not equal to -5

OpenStudy (anonymous):

@chadandtricia_2009 could you rewrite the problem in a better form.

OpenStudy (anonymous):

\[x-2divx^2+3x-10\]

OpenStudy (anonymous):

factor the denominator: = ( x -2)(x + 5)

OpenStudy (anonymous):

the x-2 ' s cancel out giving 1 / (x + 5)

OpenStudy (anonymous):

so now you should know the correct option

OpenStudy (anonymous):

so it is D

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thanks can you help me if I need anymore help?

OpenStudy (anonymous):

when x = -5 the denominator = 0 x = -5 is the excluded value

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Simplify the rational expression. State any excluded values. x^2-36/6-x A. –x – 6; where x is not equal to 6 B. x + 6; where x is not equal to -6 C. 1/x-6 ; where x is not equal to 6 D. 1/x+6 ; where x is not equal to -6

OpenStudy (anonymous):

the numerator is difference of 2 sqaures:

OpenStudy (anonymous):

I think the answr is B

OpenStudy (anonymous):

= (x - 6)(x + 6) ------------- = - (x + 6) = -x - 6 -(x - 6)

OpenStudy (anonymous):

what value of x will make the denominator zero?

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

no its 6 ; -(6 - 6) = 0

OpenStudy (anonymous):

so that elimates a and c

OpenStudy (anonymous):

so it is b or d

OpenStudy (anonymous):

nope its A

OpenStudy (anonymous):

ok thanks again

OpenStudy (anonymous):

note what i had to do in oreder to simplify i had to change 6 - x to -(x - 6) in order to cancel out the x - 6 in the numerators